OpenStudy (anonymous):

A hotel offers two activity packages. One costs $192 and includes 3 h of horseback riding and 2 h of parasailing. The second costs $213 and includes 2 h of horseback riding and 3 h of parasailing. What is the cost for 1 h of each activity?

OpenStudy (texaschic101):

3h + 2p = 192 -->(-2) 2h + 3p = 213 -->(3) ------------- -6h - 4p = - 384 (result of multiplying by -2) 6h + 9p = 639 (result of multiplying by 3) --------------add 5p = 255 p = 255/5 p = 51 now sub 51 in for p in either of the original equations to find h 3h + 2p = 192 3h + 2(51) = 192 3h + 102 = 192 3h = 192 - 102 3h = 90 h = 90/3 h = 30 check... 3h + 2p = 192 2h + 3p = 213 3(30) + 2(51) = 192 2(30) + 3(51) = 213 90 + 102 = 192 60 + 153 = 213 192 = 192 (correct) 213 = 213 (correct) 1 hr of horseback riding (h) costs $ 30 1 hr of para-sailing (p) costs $ 51

OpenStudy (texaschic101):

I will be happy to answer any questions that you may have :)

OpenStudy (anonymous):

Thank you, I'll be sure to let you know if I have anything else.

OpenStudy (texaschic101):

anytime :)